Suppose the following data describe a nation’s population
gap siven the Or found in part (o) and -6 understanding of market economies und tdhe uctois teys ned topleasc discuss factors that may cause inlarion or unenaoymcn t6gy Wagal Assuming that the inflation rate from 2015 a $50,000 income in each of the (6 pts) 4. ree years? (Show all workt) What doses t through 2017 is 6%per year, what would be the real value of s. Why are prices important in a market economy? Explain.(l p) 6. Suppose the following data describe a nations population Population Labor Force Year 1 310 million 150 million Year 2 315 million 160 million Unemployment rate 8 percen 7 pertan a. How many people are unemployed in each year? (pt) If in Year 1 the full-employment level of output was $3,00 billion and thc fll-employont rate of unemployment was 5 Okuns Law? (4 pts) hHow many people are employed in each year? (5 pts) b. .5%, how much potential GDP was lost in Year l based C. (6 different types of unemployment mentioned ida d. List and briefly discuss the 0012 what was the inflation rate between 2012 and
Answer #1

a) Number of people unemployed = Labour force * Unemployment rate

In year 1

Labour force = 150 million

Unemployment rate = 8% = 0.08

Number of people unemployed = 0.08*150 million = 12 million

In year 2

Labour force = 160 million

Unemployment rate = 7% = 0.07

Number of people unemployed = 0.07*160 million = 11.2 million

b) People employed = labour force - number of unemployed

In year 1 people employed = 150-12 = 138 million

In year 2 people employed = 160-11.2 = 148.8 million.

d) Types of unemployment :

1) Structural unemployment : When the skills of job required does not match with the worker skills. This unemployment will occur

2) Cyclical unemployment: Unemployment due to economic downturn is called cyclical unemployment

3) Frictional unemployment: When people change from one job to another the period at which they are unemployed is called frictional unemployment

4) Seasonal unemployment: When the people work only for particular season and will not work for remaining season. This is called seasonal unemployment.